# To Offload or Not To Offload: Model-driven Comparison of Edge-native and On-device Processing

### v1

@@ 3.2.1. Comparing Edge Offload and On-Device Execution
Remark 3.2 shows how the computational demand of the workload influences the execution strategy. For workloads with higher computational demand, edge offloading is more likely to outperform on-device processing, as the edge’s processing speed advantage becomes larger. Conversely, for workloads with lower computational demand, on-device processing is more efficient as the performance advantage of edge processing is likely outweighed by the associated network costs. Remark 3.3 demonstrates how network delays affect the efficiency of edge offloading. With smaller network bandwidth or larger request/result payloads, both the network queuing delay and transmission time increase. As a result, edge offloading becomes less efficient, while on-device processing is unaffected since processing is local and does not rely on network resources.
3.3. Impact of Multi-Tenancy
Figure 2. $m$ devices offloading requests to an edge server.
While the analysis in the Sec 3.2.1 implicitly assumes an edge server that is dedicated to a client device, in practice, edge servers will be multiplexed across multiple client devices and service multiple applications that offload their work. Hence, we extend our model-driven comparison to multi-tenancy scenarios.
@@ 3.3. Impact of Multi-Tenancy
Since the request arrivals at each client device are governed by independent Poisson process, and the composition of independent Poisson processes results in a new Poisson process (Pishro-Nik, 2014), it follows that the aggregate edge workload is also Poisson with an arrival rate $\lambda_{edge}$ that is the sum of the device workload. That is, $\lambda_{edge}=\sum_{i=0}^{m}\lambda_{i}.$, where $\lambda_{i}$ denotes the workload at the $i-th$ device.
We assume that the multiplexed applications have different service time demands, and the aggregate service time distribution seen at the edge will be arbitrary. For this case, we model each system as an $M/G/1$ queue, where $M$ denotes Poisson arrivals and $G$ denotes the general (arbitrary) service time distribution seen at the device and the edge. This extension captures the potentially large variability in service times across applications in multi-tenant environments. If $s_{i}$ denotes the average service time of requests of device $i$ when processed at the edge, then $s_{edge}$ is the weighted average of these service times. That is, $s_{edge}=\sum_{i}\frac{\lambda_{i}}{\lambda_{edge}}s_{i}$ and the the the effective service rate of the edge $\mu_{edge}=1/s_{edge}$. Then the aggregate utilization of the edge server $\rho_{edge}$ is $\rho_{edge}=\lambda_{edge}/\mu_{edge}^{proc}$. This leads us to the following lemma for the multi-tenant case.
Lemma 3.0.
For multi-tenant edge servers, the average end-to-end latency of edge offloading is higher than on-device processing when {dmath} s_dev - s_edge ¡ λdevμdevnetdevnet- λdev) + λedgeμedgenetedgenet- λedge) + Dreq+ DresB + ρedge+ λedgeμedgeprocVar[sedge]2(μedgeproc- λedge) - 12 ( 1kdevμdevproc- λdev - 1kdevμdevproc )
@@
The expected wait time of $M/G/1/FCFS$ systems is well-known and is given by the Pollaczek-Khinchin (P-K) formula (Gross et al., 2008; Harchol-Balter, 2013): {dmath} E[w] = ρ+ λμVar[s]2(μ- λ) Substituting (3.3) for $w_{edge}^{proc}$, (5) for $w_{dev}^{proc}$, and (6) for $w_{dev}^{net}$ and $w_{edge}^{net}$ into Equation 4 completes the proof. ∎
Practical takeaways. The expected wait time in an $M/G/1/FCFS$ system is heavily affected by the variance in service times, meaning that co-located applications with highly variable service demands can significantly increase queuing delays. It is therefore important for application designers to consider service time variability when assigning applications to edge servers. Grouping applications with similar processing demands can help reduce overall variability, thereby minimizing queuing delays and improving edge offloading performance. Further, the above analysis assumes that isolation features such as GPU virtualization, which are available on server-grade GPUs, are not available on edge accelerators. If such features were available, each device could be assigned to an isolated virtual GPU on the edge. In this special case, the offloading situation is reduced to a situation similar to that considered in Sec 3.2.1 where each virtual GPU presents $\frac{1}{m}$-th of the capacity to each device. However, many edge accelerators lack such features, resulting in a shared accelerator that services a combined workload.
4. Model Validation
@@
4.1. Experimental Setup
Workloads. We use DNN-based accelerator-driven workloads to validate our model’s ability to capture the impact of workload characteristics. We focus on cognitive assistant applications that use different DNN models for image classification and object detection. Specifically, we use three widely deployed DNN architectures for image classification and object detection, namely MobileNet, YOLO, and Inception. These models exhibit diverse computational demands and input dimensionalities as shown below:
@@ 4.1. Experimental Setup
MobileNetV2 $[1,224,224,3]$ 3.5 0.6
InceptionV4 $[1,299,299,3]$ 42.7 6.3
YoloV8n $[1,640,640,3]$ 3.2 8.7
@@
All inference tasks are executed using TensorFlow 2.13.0 across the evaluated hardware. We use frames extracted from a 720p dash cam video as input to the models, with each frame processed either locally on the device or offloaded to the edge. Since application workloads are often dynamic (e.g., triggered by motion detection), we implement a workload generator that selects video frames and generates requests following a Poisson process. The workload generator runs on a separate machine to avoid performance interference with the device during evaluation.
@@
Hardware. We use a diverse hardware platform to reflect real-world deployment scenarios. For mobile devices with accelerators, we use an NVIDIA Jetson TX2 and an NVIDIA Jetson Orin Nano, with TX2 being an earlier model and the Orin Nano a newer, more powerful model. For edge servers, we use a Dell PowerEdge R630 equipped with an NVIDIA A2 GPU, a 40-core Intel Xeon E5-2660v3 CPU, and 256GB RAM. We also use a server with an NVIDIA RTX 4070 GPU, 8-Core AMD Ryzen 7 5700X CPU, and 32GB RAM. Both edge servers are connected to a 1 Gbps network. These hardware platforms span a wide range of performance levels, from the TX2’s 1.3 TFLOPS peak FP16 performance to the 2.6 TFLOPS of the Orin Nano, the 4.5 TFLOPS of the A2, and the 29.1 TFLOPS of the RTX 4070. For each accelerator, the effective degree of parallelism $k$ depends on both the number of cores and the workload’s processing and memory demands, and is not directly observable. To estimate $k$, we empirically measure how response time varies with request rate for each workload and identify a value of $k$ that best captures the observed scaling behavior. All platforms run Ubuntu 18.04, and we utilize the Linux Traffic Control (TC) subsystem to emulate a range of network bandwidth in our experiments.
4.2. Parameter Estimation
@@
Our model requires several parameters to predict the end-to-end response time for each execution strategy. Here, we describe how these parameters are obtained in our evaluation and outline other practical methods for estimating these parameters.
Estimating average service times. We estimate request service times by monitoring inference durations using the NVIDIA System Management Interface (nvidia-smi) tool (NVIDIA, 2024), which reports per-process execution times on the accelerator. There is also extensive literature (Wang et al., 2024b; Yao et al., 2018; Kang et al., 2017; Wang et al., 2024b) on estimating the latency of neural network inference by analyzing factors such as network structure and parameter size, which achieves high accuracy due to the constant number of operations per inference across requests.
@@
Estimating average network delays. Estimating transmission delay requires knowledge of the available network bandwidth, which we continuously measure in real-time using iperf (The iPerf Project, 2025) to quickly adjust to changes. In addition to queuing and transmission delays captured by our analytic models, network delays also include propagation and processing delays. Since we assume that requests are offloaded to a nearby edge server, the propagation delay (caused by signals traveling at the speed of light) is minimal and thus excluded from our models. Processing delay can be estimated by measuring the round-trip time of small probe packets, and is excluded from our models due to its negligible impact on overall latency.
@@
4.3. Impact of Workload Characteristics
Our first experiment investigates how workload characteristics influence offloading decisions. Specifically, we evaluate three workloads with distinct computational profiles: 1) MobileNetV2 (low compute, small payload); 2) InceptionV4 (high compute, moderate payload); and 3) YOLOv8n (high compute, large payload due to its larger input dimensions). We compare on-device processing on TX2 and Orin Nano against offloading workloads to an NVIDIA A2 GPU, using a request rate of 2 requests per second (RPS) and a network speed of 5 Mbps. For each device, we test all available GPU frequency settings.
Figure 3. Latency comparison across workloads at 5 Mbps network speed and 2 RPS. (a-b) MobileNetV2, (c-d) InceptionV4, and (e-f) YOLOv8n. Left: TX2 vs A2, Right: Orin Nano vs A2.
Figure 2 compares the average latency of on-device processing across various GPU frequency settings with edge offloading under a 5 Mbps network. Figures 3(a) and 3(b) show that for MobileNetV2 (low compute, small payload), local processing on both devices outperform offloading: TX2 achieves lower latency in its three highest frequency modes, while Orin Nano consistently outperforms offloading across all frequency levels. In contrast, the trend reverses for InceptionV4 (high compute, moderate payload), where offloading provides lower average latency across all device configurations, as shown in Figures 3(c) and 3(d). Figures 3(e) and 3(f) plot the average latencies for YOLOv8n (high compute, large payload). In contrast to the results for InceptionV4, TX2 outperforms offloading in its top six performance modes, despite YOLOv8n’s high compute demand. Meanwhile, Orin Nano achieves lower average latency than offloading across all performance modes. We observe similar results when offloading to RTX4070, and omit the details here due to space constraints. In all cases, the model predictions closely match the observed response times for both approaches, achieving a mean absolute percentage error of 2.2%, with 91.5% of predictions falling within $\pm$5of the observed latency and 100% within $\pm$10%, proving the accuracy of our models.
Collectively, these empirical results align with the qualitative observations outlined in Remark 3.2 and Remark 3.3. While the performance gap between edge and device is fixed, the absolute advantage of offloading scales with the workload’s demands. For workloads with low computational demand (e.g., MobileNetV2), the edge’s relative speed advantage results in only small absolute gains that can be overshadowed by network delay. However, for computationally intensive workloads (e.g., InceptionV4), the same performance ratio yields substantially larger absolute improvements that outweigh the network costs of offloading. Moreover, workloads with large request payloads (e.g., YoloV8n) favor on-device processing because it does not involve any network components and can avoid involving a large network delay.
@@ 4.3. Impact of Workload Characteristics
4.4. Impact of Network Conditions
Figure 4. Latency comparison across execution strategies for MobileNetV2 at 2 RPS with varying network bandwidth: (a) RTX4070 as the edge server, (b) A2 as the edge server.
Figure 4 evaluates how network speed affects the effectiveness of each execution strategy. Figure LABEL:plot:mnet_varyNet_4070 compares the latency of on-device processing on TX2 and Orin Nano with edge offloading to RTX4070 under varying network speeds. For both devices, we observe a performance crossover point where offloading becomes faster than on-device processing. The location of the crossover depends on both network bandwidth and the compute performance gap between the device and the edge. For the lower-performance TX2, offloading becomes advantageous at 5 Mbps, while the faster Orin Nano requires a higher bandwidth (15 Mbps) for offloading to be beneficial.
Figure LABEL:plot:mnet_varyNet_A2 further demonstrates the relationship between network speed and device-edge performance difference. We find that even for a single device, multiple crossover points may exist depending on network conditions and the device’s performance. For example, at 5 Mbps, the device outperforms edge offloading under its highest-performance mode. However, as network bandwidth improves to 10 Mbps, edge offloading becomes more favorable. Across all scenarios, our analytical models closely predict the crossover points, demonstrating their accuracy in capturing the tradeoff between the computation speedup of edge offloading and network delay.

### v2

@@ III-C1 Comparing Edge Offload and On-Device Execution
Remark III.1 shows how the computational demand of the workload influences the execution strategy. For workloads with higher computational demand, edge offloading is more likely to outperform on-device processing, as the edge’s processing speed advantage becomes larger. Conversely, for workloads with lower computational demand, on-device processing is more efficient as the performance advantage of edge processing is likely outweighed by the associated network costs. Remark III.2 demonstrates how network delays affect the efficiency of edge offloading. With smaller network bandwidth or larger request/result payloads, both the network queuing delay and transmission time increase. As a result, edge offloading becomes less efficient, while on-device processing is unaffected since it does not rely on network resources.
Extending to collaborative processing. Collaborative (i.e., split) processing involves partially processing requests locally on the client’s device before offloading them to a remote edge server to minimize the amount of data sent over the network [34, 35, 36, 12, 26, 37, 38]. To compute the end-to-end response time of such a strategy, we combine our analytic models for on-device processing and edge offloading into a tandem queuing network that accounts for both the partial local computation and the subsequent transmission and processing delays at the edge server. Specifically, a request first enters the components modeled in Figure 1b and is partially processed with a service time $s^{\prime}_{dev}$ representing the partial local processing time. Then, the intermediate output of size $D_{inter}$ passes through the components modeled in Figure 1a, with a network transmission time of $D_{inter}/B$ and a service time $s^{\prime}_{edge}$ representing the remaining computation time on the edge. This combined model can then be used to estimate performance and inform decision-making when selecting between local, edge, or collaborative processing strategies.
III-D Impact of Multi-Tenancy
Figure 2: $m$ devices offloading requests to an edge server.
While the analysis in Lemma III.1 implicitly assumes an edge server that is dedicated to a client device, in practice, edge servers will be multiplexed across multiple client devices and service multiple applications that offload their work. Hence, we extend our model-driven comparison to multi-tenancy scenarios.
@@ III-D Impact of Multi-Tenancy
Since the request arrivals at each client device are governed by independent Poisson process, and the composition of independent Poisson processes results in a new Poisson process [39], it follows that the aggregate edge workload is also Poisson with an arrival rate $\lambda_{edge}$ that is the sum of the device workload. That is, $\lambda_{edge}=\sum_{i=0}^{m}\lambda_{i}.$, where $\lambda_{i}$ denotes the workload at the $i-th$ device.
We assume that the multiplexed applications have different service time demands, and the aggregate service time distribution seen at the edge will be arbitrary. For this case, we model each system as an $M/G/1$ system, where $M$ denotes Poisson arrivals and $G$ denotes the general (arbitrary) service time distribution seen at the device and the edge. This extension captures the potentially large variability in service times across applications in multi-tenant environments. If $s_{i}$ denotes the average service time of requests of device $i$ when processed at the edge, then $s_{edge}$ is the weighted average of these service times. That is, $s_{edge}=\sum_{i}\frac{\lambda_{i}}{\lambda_{edge}}s_{i}$ and the the the effective service rate of the edge $\mu_{edge}=1/s_{edge}$. Then the aggregate utilization of the edge server $\rho_{edge}$ is $\rho_{edge}=\lambda_{edge}/\mu_{edge}^{proc}$. This leads us to the following lemma for the multi-tenant case.
Lemma III.2.
For multi-tenant edge servers, the average end-to-end latency of edge offloading is higher than on-device processing when {dmath} s_dev - s_edge ¡ λdevμdevnetdevnet- λdev) + λedgeμedgenetedgenet- λedge) + Dreq+ DresB + ρedge+ λedgekedgeμedgeprocVar[sedge]2(kedgeμedgeproc- λedge) - 12 ( 1kdevμdevproc- λdev - 1kdevμdevproc )
@@
The expected wait time of $M/G/1/FCFS$ systems is well-known and is given by the Pollaczek-Khinchin (P-K) formula [33, 19] {dmath} E[w] = ρ+ λμVar[s]2(μ- λ) Substituting (III-D) for $w_{edge}^{proc}$, (5) for $w_{dev}^{proc}$, and (6) for $w_{dev}^{net}$ and $w_{edge}^{net}$ into (4) completes the proof. ∎
Practical takeaways. The expected wait time in $M/G/1/FCFS$ systems is heavily affected by the variance in service times, meaning that co-located applications with highly variable service demands can significantly increase queuing delays. It is therefore important for application designers to consider service time variability when assigning applications to edge servers. Grouping applications with similar processing demands can help reduce overall variability, thereby minimizing queuing delays and improving edge offloading performance. Further, the above analysis assumes that isolation features such as GPU virtualization, which are available on server-grade GPUs, are not available on edge accelerators. If such features were available, each device could be assigned to an isolated virtual GPU on the edge. In this special case, the offloading situation is reduced to a situation similar to that considered in Sec III-C1 where each virtual GPU presents $\frac{1}{m}$-th of the capacity to each device. However, many edge accelerators lack such features, resulting in a shared accelerator that services a combined workload.
III-E Discussion
Generalizing to other accelerator-driven workloads. While the above analysis focuses on DNN workloads with deterministic service times modeled using an $M/D/1$ system, our modeling framework can be extended to other workloads with different service time characteristics by incorporating appropriate queuing results. For example, workloads with recurrent architectures such as GRUs and LSTMs often process variable-length inputs with service times that vary with the input sequence length. Similarly, early-exit DNNs allow requests to exit at different layers based on confidence thresholds, which results in variable service times. To capture such variability, our framework can adopt an $M/M/1$ queuing system where service times follow an exponential distribution with the expected wait time shown in (6). By selecting the appropriate queuing model, our framework can adapt to diverse workloads and provide accurate response time estimates as will be shown in Section IV-D. Currently, our framework does not account for workloads that violate the FCFS assumption, such as Large Language Model (LLM) inference with continuous batching where requests are dynamically batched, and extending the models to capture such behavior is a subject of future work.
Model limitations. Our analysis assumes request arrivals follow a Poisson process and are independent, which may not hold under bursty arrivals. To address this, our models can adopt a $G/G/1$ system, which accommodates arbitrary arrival and service time distributions. Although no closed-form solution exists for the expected wait time in this setting, Marshall et al. [40] provide an upper bound {dmath} E[w_G/G/1] ≤λ(σa2+ σs2)2(1-ρ) where $\sigma_{a}$ and $\sigma_{s}$ are the variances for the interarrival and service distributions respectively. Additionally, the queuing results used in our analysis assume that each hardware queue has infinite capacity. For scenarios with bounded queue sizes, we can incorporate results for finite-buffer queuing systems [19] to model the impact of such capacity constraints.
IV Model Validation
@@
IV-A Experimental Setup
Workloads. We validate our model’s ability to capture the impact of workload characteristics using a diverse set of workloads, including DNN-based models as well as recurrent models. For DNN workloads, we focus on cognitive assistant applications involving image classification and object detection. Specifically, we use three widely deployed DNN architectures, namely MobileNet, Inception, and YOLO, each with distinct computational demands and input sizes as shown below:
@@ IV-A Experimental Setup
MobileNetV2 $[224,224,3]$ 3.5 0.6
InceptionV4 $[299,299,3]$ 42.7 6.3
YOLOv8n $[640,640,3]$ 3.2 8.7
@@
All inference requests are executed using TensorFlow 2.13.0 across the evaluated hardware. We use frames extracted from a 720p dash cam video as input to the models, with each frame processed either locally on the device or offloaded to the edge. Since application workloads are often dynamic (e.g., triggered by motion detection), we implement a workload generator that selects video frames and generates requests following a Poisson process, which runs on a separate machine to avoid performance interference with the device during evaluation. For recurrent models, we focus on personal voice assistant applications involving sentiment classification of conversational utterances. We validate our models using two types of recurrent neural networks: a GRU model with lower computational demand and a more compute-intensive LSTM model, with their architectural specifications summarized below:
Model Embedding Dim # Layers $\times$ Units Activation
GRU 512 2 $\times$ 512 Sigmoid
LSTM 1024 3 $\times$ 1024 Sigmoid
Both models are implemented using TensorFlow’s Keras API. To simulate real-world user inputs, we use the Amazon Alexa Topical Chat dataset [41], a knowledge-grounded human-to-human conversation dataset consisting of spoken-style dialogues with an average utterance length of 19 tokens.
@@
Hardware. We use a diverse set of hardware platforms to reflect real-world deployment scenarios. For client devices with accelerators, we use the NVIDIA Jetson TX2 and Jetson Orin Nano, two commonly deployed platforms in IoT applications, with the TX2 being an earlier model and the Orin Nano offering better performance. For edge servers, we use a Dell PowerEdge R630 equipped with an NVIDIA A2 GPU, a 40-core Intel Xeon E5-2660v3 CPU, and 256GB RAM. We also use a server with an NVIDIA RTX 4070 GPU, 8-Core AMD Ryzen 7 5700X CPU, and 32GB RAM. Both edge servers are connected to a 1 Gbps network. These platforms span a wide performance range, from 1.3 TFLOPS (TX2) to 2.6 TFLOPS (Orin Nano), 4.5 TFLOPS (A2), and up to 29.1 TFLOPS (RTX 4070) in peak FP16 throughput. For each accelerator, the effective degree of parallelism $k$ depends on both the number of cores and the workload’s processing and memory demands, and is not directly observable. To estimate $k$, we empirically measure how response time varies with request rate for each workload and identify a value of $k$ that best captures the observed scaling behavior. All platforms run Ubuntu 18.04, and we use the Linux Traffic Control (TC) subsystem to emulate different network bandwidths.
IV-B Parameter Estimation
@@
Our model requires several parameters to compute the end-to-end response times. We now describe how these parameters are obtained in our evaluation and outline other practical methods for estimating them.
Estimating average service times. Our system estimates workload service times through empirical profiling and a neural network–based prediction approach. We estimate service times for DNN workloads by monitoring inference durations using the NVIDIA System Management Interface (nvidia-smi) tool [23], which reports per-process execution times on the accelerator. For GRU and LSTM workloads, we profile request latencies on a representative input set distinct from the test set and use the average latency from this profiling as the service time of the model. This profiling data is used for our edge-offloading and on-device processing experiments. For collaborative processing, a neural network is trained using profiling data to predict service time based on prior work discussed in Section III-B. Such a neural network approach avoids the extensive profiling overhead of all possible split configurations by predicting partial execution service times.
@@
Estimating average network delays. Estimating transmission delay requires knowledge of the available network bandwidth, which we continuously measure in real-time using iperf [42]. In addition to queuing and transmission delays captured by our models, network delays also include propagation and processing delays. Since we assume that requests are offloaded to a nearby edge server, the propagation delay (caused by signals traveling at the speed of light) is minimal and thus excluded from our models. Processing delay can be estimated by measuring the round-trip time of small probe packets, and is excluded from our models due to its negligible impact on overall latency.
@@
IV-C Impact of Workload Characteristics
Our first experiment investigates how workload characteristics influence offloading decisions using three DNN workloads with distinct profiles: 1) MobileNetV2 (low compute, small payload); 2) InceptionV4 (high compute, moderate payload); and 3) YOLOv8n (high compute, large payload). We compare on-device processing on TX2 and Orin Nano against offloading to an A2 GPU using a request rate of 2 requests per second (RPS) and a network speed of 5 Mbps. For each device, we test all available GPU frequency settings to emulate different device conditions.
Figure 3: Latency comparison across workloads at 5 Mbps network speed and 2 RPS. (a-b) MobileNetV2, (c-d) InceptionV4, and (e-f) YOLOv8n. Left: TX2 vs A2; Right: Orin Nano vs A2.
Figure 3 compares the average latency of on-device processing and edge offloading across workloads, devices, and GPU frequency levels. Figures 3a and 3b show that for MobileNetV2, local processing on both devices outperform offloading: TX2 achieves lower latency in its three highest frequency modes, while Orin Nano consistently outperforms offloading across all frequency levels. The trend reverses for InceptionV4 as shown in Figures 3c and 3d, where offloading provides lower average latency across all device configurations. Lastly, Figures 3e and 3f shows the results for YOLOv8n. In contrast to the results for InceptionV4, TX2 outperforms offloading in its top six performance modes despite YOLOv8n’s high compute demand, while Orin Nano achieves lower average latency than offloading across all performance modes. We observe similar results when offloading to RTX4070 and omit the details here due to space constraints. In all cases, the model predictions closely match the observed response times for both approaches, achieving a mean absolute percentage error of 2.2%, with 91.5% of predictions falling within $\pm$5of the observed latency and 100% within $\pm$10%, proving the accuracy of our models.
Collectively, these empirical results align with the qualitative observations stated in Remark III.1 and Remark III.2. While the performance ratio between edge and device is fixed, the absolute advantage of offloading scales with the workload’s demands. For workloads with low computational demand (MobileNetV2), the edge’s relative speed advantage results in only small absolute gains that can be overshadowed by network delays of offloading. However, for compute-intensive workloads (InceptionV4), the same performance ratio yields substantially larger absolute improvements that outweigh the network delays. Moreover, workloads with large message payloads (YOLOv8n) favor on-device processing because it does not involve any network components and thus avoids any network delay.
@@ IV-C Impact of Workload Characteristics
IV-D Impact of Complex Deep Learning Models
Figure 4: Latency comparison across execution strategies for (a) GRU model and (b) LSTM model at 3 RPS under 10 Mbps.
Figure 5: Latency comparison across execution strategies for MobileNetV2 at 2 RPS with varying network bandwidth: (a) RTX4070 as the edge server, (b) A2 as the edge server.
Figure 6: Latency comparisons across (a) different collaborative processing configurations, (b) request rates under 10 Mbps and 20 Mbps networks, and (c) numbers of co-located applications.
In addition to the DNN workloads, our model also accurately captures the latency behavior of more complex deep learning workloads. Figure 4 compares the average latency of on-device processing on the Orin Nano against offloading to A2 for two recurrent models: GRU and LSTM, under a moderate load of 3 RPS. Since service times vary with input length in these models, we incorporate the $M/M/1$ queuing model into our analysis to account for this variability in latency prediction.
Figure 4a shows that even for the lightweight GRU model, offloading offers lower latency due to the small input utterance length (on average less than 100 bytes), which incurs minimal network delay. Figure 4b shows that as the model complexity increases, the absolute performance gain of offloading increases, leading to a greater latency reduction. In both cases, our model predictions closely match the observed latencies.
Key takeaway. Our models generalize to complex deep learning workloads beyond DNN inference by integrating appropriate queuing models.
IV-E Impact of Network Conditions
Figure 5 shows how network speed affects the effectiveness of each execution strategy. Figure 5a compares the latency of on-device processing on TX2 and Orin Nano with edge offloading to RTX4070 under varying network speeds. For both devices, we observe a performance crossover point where offloading becomes faster than on-device processing. The location of the crossover depends on both network bandwidth and the performance gap between the device and the edge. For the lower-performance TX2, offloading becomes advantageous at 5 Mbps, while the faster Orin Nano requires a higher bandwidth (15 Mbps) for offloading to be beneficial.
Figure 5b shows that even for a single device, multiple crossover points may exist depending on network conditions and the device’s performance. For example, at 5 Mbps, TX2 processing in its top three performance modes outperforms offloading; however, as TX2’s performance decreases beyond those modes, offloading becomes more advantageous. Across all scenarios, our models closely predict the crossover points, demonstrating their accuracy in capturing the tradeoff between the computation speedup of edge offloading and network delay.